OpenStudy (anonymous):

Which of the following is an example of the inverse of an exponential function? y = x/4 y = 4x y = 2^x y = log(subscript 2)x

OpenStudy (anonymous):

The inverse of an exponential function is a logarithm. Logarithms, are, essentially, "finding" the exponent. So the correct answer is the last one - it's the only logarithmic function.

OpenStudy (anonymous):

Thanks so much. I'm not so proficient in math and I was struggling on this one. Are you good at Algebra II? If so, can I ask you more questions?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Okay give me a second.

OpenStudy (anonymous):

What is the range of this? I can't figure it out for the life of me/

OpenStudy (anonymous):

Try plugging in a negative number, say, -2. You'd get y = (1/2)^-2 + 3. Since it's a negative exponent, it becomes a fraction and you'd get y = 2^2 + 3. The pattern would continue with negative numbers. And when you plug in any positive number, the fraction becomes smaller, but never really reaches 0. If you plug in 0, you'd get the least possible number. You'd get 3. You can never get less than 3. Therefore, the range would be y >= 3.

OpenStudy (anonymous):

The best way to deal with range problems is to plug in 0, a negative, and a positive number.

OpenStudy (anonymous):

Thank you so much. Do you mind if I ask you a few more that I'm stuck on?

OpenStudy (anonymous):

Not at all

OpenStudy (anonymous):

Thank you so much. This one is giving me trouble. Every answer I get is not one of my choices. Write log(subscript 5) 625 = 4 in exponential form.

OpenStudy (anonymous):

The subscript is base. So it's read log base 5. Since 5 is the base, you then raise it to the power of whatever the logarithm is equal to - in this case 4. Then you set that equal to what's in the middle. The way I remember it is start from the left, go to the right, then the middle. In exponential form it's 5^4 = 625.

OpenStudy (anonymous):

Ah thank you so much. You're awesome.

OpenStudy (anonymous):

No problem. Any more?

OpenStudy (anonymous):

I might have a few, yeah. But I actually have to get going. Can I message you for help?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Fantastic. Thank you so much.